# Automated Disentangling Analysis of Skin Colour for Lesion Images

### v1

@@ Abstract
Machine-learning models working on skin images often have degraded performance when the skin colour captured in images (SCCI) differs between training and deployment. Such differences arise from entangled environmental factors (e.g., illumination, camera settings), and intrinsic factors (e.g., skin tone) that cannot be accurately described by a single “skin tone” scalar. To mitigate such colour mismatch, we propose a skin-colour disentangling framework that adapts disentanglement-by-compression to learn a structured, manipulable latent space for SCCI from unlabelled dermatology images. To prevent information leakage that hinders proper learning of dark colour features, we introduce a randomized, mostly monotonic decolourization mapping. To suppress unintended colour shifts of localized patterns (e.g., ink marks, scars) during colour manipulation, we further propose a geometry-aligned post-processing step. Together, these components enable faithful counterfactual editing and answering an essential question: "What would this skin condition look like under a different SCCI?", as well as direct colour transfer between images and controlled traversal along physically meaningful directions (e.g., blood perfusion, camera white balance), enabling educational visualization of skin conditions under varying SCCI. We demonstrate that dataset-level augmentation and colour normalization based on our framework achieve competitive lesion classification performance.
@@ 1 Introduction
Existing mitigation strategies for colour mismatch fall into two categories. First, most recent works in dermatology focus on augmenting the training set using augmented images [1] or partially or fully synthetic images [16, 8]. While these methods show promising performance improvement, they only focus on intrinsic skin-tone change and usually consider skin colour on individual images without a structured colour model, offering limited control of the result and distribution matching of datasets. The single-value scalar or categorical skin-tone descriptors (such as FST [7] and MST [15]) used by these methods have been shown to have very poor correlation with the skin colour shown in photos [10, 23], because it does not reflect other factors (including blood perfusion and environmental factors). Second, colour normalization aims at mitigating the mismatch by converting each image to a “standard” colour. While such a method has been widely used in histology [4, 24, 22], only limited works have been proposed for skin images [12, 20, 19], possibly due to more complex colour formation process. Currently, these methods focus on environmental factors without explicitly considering the intrinsic factors that affect the SCCI. Moreover, educators and future physicians need a tool for visualizing the same skin condition under different SCCI, which requires a representation that can be traversed in a controlled and interpretable way.
@@
In this work, we propose an unsupervisedly learned skin-colour disentangling framework that treats SCCI as the outcome of multiple entangled factors. Enabled by recent applications of the information bottleneck principle [21] to low-level vision feature disentanglement [13, 25], our method adapts [25]’s disentanglement-by-compression framework and compresses skin colour into an organized latent space whose entries are approximately independent and can be individually adjusted. Faithful disentanglement of SCCI possesses unique challenges compared to [25]. First, to avoid shortcuts that leak skin-darkness/shading information and obstruct the representation learning, we introduce a randomized, a mostly monotonic decolourization mapping that generates the colourless input required by the framework. Second, to suppress the unintended colour shift of localized patterns (e.g., ink marks, scars), we propose a geometry-aligned post-processing step to selectively reject unreliable changes. The ability of faithful controlled manipulation in a latent space enables visualizing the lesion under a different SCCI.
@@
After training, SCCI can be transferred directly from source images to targets (fig.˜1(a)), individual latent entries can be adjusted independently (fig.˜1(b)), dataset-level augmentation can be performed by sampling from a target distribution (fig.˜1(c)), and data normalization can be performed by setting the colour embedding to a standard value. Physically meaningful trajectories can be identified and traversed in the latent space (fig.˜2) for educational visualization. In summary, the main contributions of this work are:
@@ 3 Experiments and Results
Each model is trained on an H100 GPU for 50 epochs using the same loss functions, training procedure, and optimization settings as in [25], except that we set $\lambda_{\text{bpp\_g}}=0.003$, $\lambda_{\text{diver}}=0.05$, and $\lambda_{\text{color}}=0.1$. Different downstream tasks may use different training sets. For qualitative experiments, we use the augmentation model trained on the training and test sets in  [16] and ISIC-2020 [18] (both without labels). For post-processing, we set $\tau=0.1$ and $h(d)=d^{2}$. The code will be released upon publication.
@@ 3.3 Dataset Augmentation and Colour Normalization for Lesion Classification
Our method is a general framework applicable to various downstream tasks. Due to space constraints, we present only the results of training a lesion malignancy classification model. Note that the goal of these experiments is not to propose a state-of-the-art classifier, but to explore the potential of our augmentation and normalization methods and to compare their effectiveness with existing approaches. We follow the same evaluation protocol as in [16]. For fair comparison, the training set (parts of DermNet NZ [5] and ISIC-2018 JID editorial images [11]) and test set (parts of Dermatology Atlas [14]) are obtained directly from the authors of [16]. For each augmented or normalized training set, we train a ResNet-50 [9] model for lesion malignancy classification using Adam with a learning rate of $10^{-3}$, repeating each experiment ten times and reporting the mean and standard deviation of accuracy. The colour model for data augmentation is trained on unlabelled training and test sets for classification, while the colour normalization model is trained on the training set only (to improve representation quality, unlabelled ISIC-2020 [18] is added to the set). Rows 3 and 4 in table˜1 show the performance of models trained on the augmented and normalized training sets, respectively. Both outperform the baseline (row 1) and the prior style-transfer augmentation method  [16] (row 2).
@@ 3.4 Ablation Study
Effects of Latent Sampling Methods Under ideal conditions (a very large training set and a highly flexible model), the entries of $\bm{e}$ may be assumed independent; in practice, however, weak correlations between entries are observed. We therefore compare several latent sampling strategies for the augmentation method; results are shown at the bottom of table˜1. We tested 3 methods for sampling the latent vector $\bm{e}$ from the distribution of the test set: (row 4) directly reusing latents from the test set (by random selection), (row 5) fitting a normalizing flow model [6], and (row 6) sampling each entry independently from the empirical marginal distribution. All methods with a relatively accurate modelling of the marginal distribution of the latent vector can achieve comparable performance, while direct random sampling performs the best.
@@ 4 Conclusion
In this work, we presented a disentanglement-by-compression framework that learns a structured, manipulable latent representation of skin colour captured in images (SCCI) directly from unlabelled dermatology data. A randomized decolourization mapping and a geometry-aligned post-processing step were proposed to improve faithfulness of the colour manipulation; ablation studies confirmed that both contribute measurably to downstream performance and perceptual fidelity. The learned latent space supports colour transfer, independent entry manipulation, and traversal along physically meaningful trajectories (e.g., blood perfusion, camera white balance), enabling training-set augmentation, colour normalization, and educational visualization of skin conditions under varying SCCI. On a lesion malignancy classification benchmark, both augmentation and normalization strategies outperformed the unaugmented baseline and a prior style-transfer method. Future work includes applying the framework to multi-centre other downstream tasks such as segmentation.

### v2

@@ Abstract
Machine-learning models applied to skin images often have degraded performance when the skin colour captured in images (SCCI) differs between training and deployment. These discrepancies arise from a combination of entangled environmental factors (e.g., illumination, camera settings) and intrinsic factors (e.g., skin tone) that cannot be accurately described by a single “skin tone” scalar – a simplification commonly adopted by prior work. To mitigate such colour mismatches, we propose a skin-colour disentangling framework that adapts disentanglement-by-compression to learn a structured, manipulable latent space for SCCI from unlabelled dermatology images. To prevent information leakage that hinders proper learning of dark colour features, we introduce a randomized, mostly monotonic decolourization mapping. To suppress unintended colour shifts of localized patterns (e.g., ink marks, scars) during colour manipulation, we further propose a geometry-aligned post-processing step. Together, these components enable faithful counterfactual editing and answering an essential question: “What would this skin condition look like under a different SCCI?”, as well as direct colour transfer between images and controlled traversal along physically meaningful directions (e.g., blood perfusion, camera white balance), enabling educational visualization of skin conditions under varying SCCI. We demonstrate that dataset-level augmentation and colour normalization based on our framework achieve competitive lesion classification performance. Ultimately, our work promotes equitable diagnosis through creating diverse training datasets that include different skin tones and image-capturing conditions.
@@ 1 Introduction
Existing mitigation strategies for colour mismatch fall into two categories. First, most recent works in dermatology focus on augmenting the training set using augmented images [1] or partially or fully synthetic images [16, 8]. While these methods show promising performance improvement, they only focus on intrinsic skin-tone change and usually consider skin colour on individual images without a structured colour model, offering limited control of the result and distribution matching of datasets. The single-value scalar or categorical skin-tone descriptors (such as FST [7] and MST [15]) used by these methods have been shown to have very poor correlation with the skin colour shown in photos [10, 23], because it does not reflect other factors (including blood perfusion and environmental factors). Second, colour normalization aims at mitigating the mismatch by converting each image to a “standard” colour. While such a method has been widely used in histology [4, 24, 22], only limited works have been proposed for skin images [12, 20, 19], possibly due to the more complex colour formation process. Currently, these methods focus on environmental factors without explicitly considering the intrinsic factors that affect the SCCI. Moreover, educators and future physicians need a tool for visualizing the same skin condition under different SCCI, which requires a representation that can be traversed in a controlled and interpretable way.
@@
In this work, we propose an unsupervisedly learned skin-colour disentangling framework that treats SCCI as the outcome of multiple entangled factors. Enabled by recent applications of the information bottleneck principle [21] to low-level vision feature disentanglement [13, 25], our method adapts [25]’s disentanglement-by-compression framework and compresses skin colour into an organized latent space whose entries are approximately independent and can be individually adjusted. Faithful disentanglement of SCCI possesses unique challenges compared to [25]. First, to avoid shortcuts that leak skin-darkness/shading information and obstruct the representation learning, we introduce a randomized, mostly monotonic decolourization mapping that generates the colourless input required by the framework. Second, to suppress the unintended colour shift of localized patterns (e.g., ink marks, scars), we propose a geometry-aligned post-processing step to selectively reject unreliable changes. The ability of faithful controlled manipulation in a latent space enables visualizing the lesion under a different SCCI.
@@
After training, SCCI can be transferred directly from source images to targets (fig.˜1(a)), and individual latent entries can be adjusted independently (fig.˜1(b)). The proposed method improves health equity by adjusting the training set for different populations. Dataset-level augmentation can be performed by sampling from a target distribution (fig.˜1(c)), and data normalization can be performed by setting the colour embedding to a standard value. Physically meaningful trajectories can be identified and traversed in the latent space (fig.˜2) for educational visualization. In summary, the main contributions of this work are:
@@ 3 Experiments and Results
Each model is trained on an H100 GPU for 50 epochs using the same loss functions, training procedure, and optimization settings as in [25], except that we set $\lambda_{\text{bpp\_g}}=0.003$, $\lambda_{\text{diver}}=0.05$, and $\lambda_{\text{color}}=0.1$. Different downstream tasks may use different training sets. For qualitative experiments, we use the augmentation model trained on the training and test sets in  [16] and ISIC-2020 [18] (both without labels). For post-processing, we set $\tau=0.1$ and $h(d)=d^{2}$. The code will be released upon acceptance.
@@ 3.3 Dataset Augmentation and Colour Normalization for Lesion Classification
Our method is a general framework applicable to various downstream tasks. Due to space constraints, we present only the results of training a lesion malignancy classification model. Note that the goal of these experiments is not to propose a state-of-the-art classifier, but to explore the potential of our augmentation and normalization methods. For fair comparison, we followed the same evaluation protocol as in [16] and obtained the training set (parts of DermNet NZ [5] and ISIC-2018 JID editorial images [11]) and test set (parts of Dermatology Atlas [14]) directly from the authors. For each augmented or normalized training set, we train a ResNet-50 [9] model for lesion malignancy classification using Adam with a learning rate of $10^{-3}$, repeating each experiment ten times and reporting the mean and standard deviation of accuracy. The colour model for data augmentation is trained on unlabelled training and test sets for classification, while the colour normalization model is trained on the training set only (to improve representation quality, unlabelled ISIC-2020 [18] is added to the set). The models trained on the augmented (row 3 in table˜1) and normalized (row 4) sets outperform the baseline (row 1) by a large margin. They also perform favorably compared to the prior augmentation method  [16], while providing significantly more interpretability in the representation space.
@@ 3.4 Ablation Study
Effects of Latent Sampling Methods Under ideal conditions (a very large training set and a highly flexible model), the entries of $\bm{e}$ may be assumed independent; in practice, however, weak correlations between entries are observed. We therefore compare several latent sampling strategies for the augmentation method; results are shown at the bottom of table˜1. We tested 3 methods for sampling the latent vector $\bm{e}$ from the distribution of the test set: (row 3) directly reusing latents from the test set (by random selection), (row 5) fitting a normalizing flow model [6], and (row 6) sampling each entry independently from the empirical marginal distribution. All methods with a relatively accurate modelling of the marginal distribution of the latent vector can achieve comparable performance, while direct random sampling performs the best.
@@ 4 Conclusion
In this work, we presented a disentanglement-by-compression framework that learns a structured, manipulable latent representation of skin colour captured in images (SCCI) from unlabelled dermatology data. A randomized decolourization mapping and a geometry-aligned post-processing step were proposed to improve faithfulness of the colour manipulation; ablation studies confirmed that both contribute measurably to downstream performance and perceptual fidelity. The learned latent space supports colour transfer, independent entry manipulation, and traversal along physically meaningful trajectories (e.g., blood perfusion, camera white balance), enabling training set augmentation, colour normalization, and educational visualization of skin conditions under varying SCCI. On a downstream benchmark, both augmentation and normalization strategies perform favourably against prior methods. This work is a step towards achieving equitable diagnosis, where our models can be reliably adapted across different clinical settings and populations.